OpenStudy (anonymous):

Ok so from what I know you begin to solve: [(x^+2x)/(12x+54)]-[(3-x)/(8x+36)] by distributing the negative : [(-3+x)/(8x+36)] and then factoring both the denominators 8x+35=4(2x+9) and 12x+54=6(2x+9) so no i have: [(x^2+2x)/(6(2x+9))]+[(-3+x)/(4(2x+9))] now the one on the left is missing 4 and the one on the right is missing 6 right so you multiply and get: [(4x^2+8x)/(24(2x+9))]+[(6x-18)/(24(2x+9))] add the numerators because they share a denominator: (4x^2+14x-18)/(24(2x+9)) lcm= 2 so you factor 2(2x^2+7x-9)/24(2x+9) simplify.. (2x^2+7x-9)/12(2x+9) ok so what h

OpenStudy (anonymous):

wolframalpha.com (it answers these problems) says the answers (x-1)/(12)

OpenStudy (turingtest):

did you try factoring (2x^2+7x-9)

OpenStudy (turingtest):

(2x^2+7x-9) you know that you want something like (x+a)(x+b) now when you foil that out notice that we need to wind up with a coefficient of 2 on the x^2 term, so that means one of our factors will have a 2x in it (since the only factors of 2 are 2 and 1) so it should look like (2x^2+7x-9) (2x+a)(x+b) now we need to find a and b again if we foil this out it should be clear that we want ab=-9, so what are the factors of -9 ? try those for a and b and see which gets the correct cross term of 7

OpenStudy (turingtest):

(2x^2+7x-9) (2x+a)(x+b) foiling this out we get 2x^2+(a+2b)x+ab we need a+2b=7 and ab=-9 for the correct factorization, so find the correct a and b such that that works

OpenStudy (turingtest):

it basically comes down to solving a+2b=7 ab=-9

OpenStudy (turingtest):

remember we want the factors of -9, not 9

OpenStudy (turingtest):

a and/or b could be negative, so don't worry about the + sign

OpenStudy (turingtest):

the factorisation of (2x^2+7x-9) will be (2x+a)(x+b) and we need to find a and b, agreed?

OpenStudy (turingtest):

(2x+a)(x+b)=2x^2+ax+2bx+ab=2x^2+(a+2b)x+ab so a+2b=7 ab=-9 let's go down your list of factorizations of -9 the first one you have is -1 and 9, so try a=-1 and b=9, or a=9 and b=-1 and see of that gets the right cross term of 7x

OpenStudy (turingtest):

another way to do it would be to just solve the system a+2b=7 ab=-9

OpenStudy (turingtest):

you chose a=1 and b=3, but 1*3 is not -9 you need to chose a and b that are factors of -9

OpenStudy (turingtest):

you made a list, try those values... 1,-9 3,-3 -1,9 those are your three options

OpenStudy (turingtest):

you cannot choose 3 and one because that does not multiply to -9 they *must* multiply to -9 here are your options -1,9 3,-3 1,-9 so there are six possibilities: a=-1,b=9 a=3,b=-3 a=1,b=-9 or the other way around b=-1,a=9 b=3,a=-3 b=1,a=-9 try them until you find which one works (with time you will be able to do it faster in your head)

OpenStudy (anonymous):

oh ok sorry, i see now, (9)+2(-1)=7,(9)(-1)=-9

OpenStudy (turingtest):

yes :) so the factorization is...?

OpenStudy (anonymous):

(2x+9)(x-1)/12(2x+9)=(x-1)/(12)

OpenStudy (turingtest):

quite welcome :)